# Total cross section of the process $e^{+}+e^{-}\to\Sigma^{+}+\bar{\Sigma}^{-}$ including the $D$-meson loop and three-gluon contributions

### v1

@@ IV The D-meson loop mechanism
$S_{D}(s)=\frac{\alpha g_{e}\beta G(s)}{48\pi^{2}s}\int\frac{dk}{i\pi^{2}}\frac{SpD(s,k^{2})}{(k^{2}-M_{\Xi}^{2})((k-q_{+})^{2}-M_{D}^{2})((k+q_{-})^{2}-M_{D}^{2})}\times$
@@
$Z_{D}(s)=\frac{1}{s}\int\frac{dk}{i\pi^{2}}\frac{SpD(s,k^{2})}{(k^{2}-M_{\Xi}^{2})((k-q_{+})^{2}-M_{D}^{2})((k+q_{-})^{2}-M_{D}^{2})}\times$
@@
where $SpD(s,k^{2})$ is the trace of the $\gamma$-matrices over the baryon line have the following form:
$SpD(s,k^{2})=Sp[(\hat{q_{+}}+M_{\Sigma})\gamma_{5}(\hat{k}+M_{\Xi})\gamma_{5}(\hat{q_{-}}-M_{\Sigma})(\hat{k}-M_{\Sigma})]=$
@@
Thus, taking into account these two $\delta$-functions (LABEL:delta) in (31), we get the imaginary part of this quantity from $Z_{D}(s)$,
$2i\,\mbox{Im}\,Z_{D}(s)=\frac{(-2\pi i)^{2}}{s}\int\frac{dk}{i\pi^{2}}\frac{SpD(s,k^{2})}{k^{2}-M_{\Xi}^{2}}G_{\psi D\bar{D}}(s,(k+q_{-})^{2},(k-q_{+})^{2})\times$
@@
×$\quad\times SpD(s,-k_{(i)}^{2})\leavevmode\nobreak\ G_{\Xi D\Sigma}^{2}(-k_{(i)}^{2},M_{D}^{2}),\qquad s>4M_{D}^{2},$ (35)
@@
Now we want to consider the function $G_{\Xi D\Sigma}(k^{2},p^{2})$ from (31). Since $k^{2}<0$, in the $t$-channel the only dependence in the imaginary part of $Z_{D}$ is the off-mass shell of the baryon $\Xi$.
We would like to note that in Ahmadov ; Bystritskiy ; BA we used a constant for the $\Lambda DP$-vertex based on the results of Reinders ; Navarra . But, in this paper for the we use $\Xi D\Sigma$-vertex a form of the constant that corresponds to the results of Reinders ; Choe .
@@ V The three-gluon mechanism
$Z_{3g}(s)=\frac{4}{\pi^{5}s}\int\frac{dk_{1}}{k_{1}^{2}}\frac{dk_{2}}{k_{2}^{2}}\frac{dk_{3}}{k_{3}^{2}}\frac{Sp3g\leavevmode\nobreak\ \delta(q-k_{1}-k_{2}-k_{3})}{((q_{+}-k_{1})^{2}-M_{\Sigma}^{2})((q_{-}-k_{3})^{2}-M_{\Sigma}^{2})},$ (44)
where $g_{col}$ is the color factor, $g_{col}(1/4)=\left<\Sigma|d^{ijk}\leavevmode\nobreak\ T^{i}T^{j}T^{k}|\Sigma\right>=15/2$. The $Sp3g$ in (44) is the product of traces over the lines of the $\Sigma$-hyperon and $c$-quark,
$Sp3g=\mbox{Sp}[\hat{Q}_{\alpha\beta\gamma}(\hat{k}_{c}+m_{c})\gamma^{\mu}(\hat{k}_{\bar{c}}-m_{c})]\cdot\mbox{Sp}\left[(\hat{q}_{+}+M_{\Sigma})\gamma^{\alpha}(\hat{q}_{+}-\hat{k}_{1}+M_{\Sigma})\gamma^{\beta}\right.\times$
@@
For the correct normalization of this wave function, it is the factor $1/\sqrt{3}$ that provides the correct normalization of this state by 1.
@@ VI The Numerical results
For $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ in the Born approximation, in Fig. 3 we show the dependence of the total cross as a function of the center-of-mass energy $\sqrt{s}$. It is seen that at the energy of $\sqrt{s}$ = 3.4 GeV, the total cross section has a maximum value and then, with increasing energy, the total cross section smoothly decreases. A comparison with the experimental data shows that our result agrees well with the experimental points. In Fig. 3, we constructed a plot according to the theoretical results for the total cross section in the Born approximation. Our theoretical result is compared with the experimental data from the BESIII.
It should be noted that the quantities $Z_{D}(s)$ from (31) and $Z_{3g}\left(s\right)$ from (44) are considered the main parts of the general cross section, which give the corresponding contributions ($D$-meson loop and three-gluons). In Figure 6 we present the dependence of the quantity of $Z_{D}(s)$ on the total energy $\sqrt{s}$ in the range from the reaction threshold $\sqrt{s}=2M_{\Sigma}^{+}$ to $4.951\leavevmode\nobreak$GeV. It can be seen from the figure that in the real and imaginary parts of $Z_{D}(s)$ the quantity remains practically the same as in the case of the production of final states $p\bar{p}$ [Fig. 7(a) in Bystritskiy ], and $\Lambda\bar{\Lambda}$ [Fig. 6 in BA ], and $\Sigma^{0}\bar{\Sigma}^{0}$ [Fig. 6 in AAhmadov ]. Now in Fig. 7 we show the dependence of the real and imaginary parts of the quantity $Z_{3g}(s)$ on the total energy $\sqrt{s}$ in the range from the reaction threshold $\sqrt{s}=2M_{\Sigma}^{+}$ to $4.951\leavevmode\nobreak$GeV. We would like to note that in this work the general behavior of the curves of the quantity $Z_{3g}(s)$ by $\sqrt{s}$ is very different from the curves, which were produced in the final state $p\bar{p}$ [Fig.  7 (B) Bystritskiy ], $\Lambda\bar{\Lambda}$ [Fig. 7 BA ] and $\Sigma^{0}\bar{\Sigma}^{0}$ [Fig. 7 in AAhmadov ]. It can be seen from this graph that the characteristic large negative values of the quantity $Z_{3g}(s)$ remain the same, which gives a large relative phase with respect to the Born contribution to the amplitude. The resonance position $\psi(3770)$ in Figures 6 and 7 is marked by a vertical dashed line.
@@
The contributions of the pure $D$-meson loop to the Born cross section are shown in Fig.8. In Figure 9 we illustrate the result of a pure three-gluon contribution. The theoretical results obtained for both of these contributions are compared with the data of the BESIIII26 collaboration.
@@
In Fig. 10 we present the results of the dependence of the total cross section in the energy range $\sqrt{s}$ = 3.4 - 5.0 GeV taking into account the $D$-loop and three-gluon mechanisms in the Born cross section and compare it with the BESIII I26 data. In this graph, we want to show that the peak is near the resonance of the $\psi(3770)$ charmonia and is clearly visible. It is necessary to note that in our model, we mainly consider the vicinity of $\psi(3770)$. But, in addition to this, we consider the vicinity with new charmonium-like states in the same energy region, for example, $\psi(4040)$, $\psi(4160)$, $Y(4230)$, $Y(4360)$, $\psi(4415)$ and $Y(4660)$. Thus, we basically have the $\psi(3770)$ charmonium in the intermediate state, but in addition we calculate the total cross section taking into account the new charmonium-like states in the intermediate state.
@@
Now in Figure 11 we illustrate the dependence of the total relative phase $\phi_{\psi}$ from the center of mass energy $\sqrt{s}$. We determine the total relative phase $\phi_{\psi}$ as the contribution of the charmonium $\mathcal{M}_{\psi}$ to the amplitude with respect to the Born contribution of $\mathcal{M}_{B}$ without taking into account the Breit-Wigner factor, that is,
@@
In this present work, in addition to the contribution of the $\psi(3770)$ charmonium in the intermediate state, we also considered the contributions from charmonium-like states, $\psi(4040)$,   $\psi(4160)$,   $Y(4230)$,  $Y(4360)$,  $\psi(4415)$, and $Y(4660)$ in the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ process.
In Fig. 12 we present the total cross section at the center-of-mass of energy in the region $\sqrt{s}$ = 3.510 - 4.951 GeV according to the results of the contributions of charmonium (-like) states, $\psi(4040)$,   $\psi(4160)$,   $Y(4230)$,  $Y(4360)$,  $\psi(4415)$, and $Y(4660)$ in the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ process, respectively. We would like to note that in Fig. 12 it is seen that the experimental result at the point for each resonance of charmonia $\psi(4040)$,   $\psi(4160)$,   $Y(4230)$,  $Y(4360)$,  $\psi(4415)$, and $Y(4660)$ are in very good agreement with our theoretical result.
@@ VII Summary and Conclusion
In the present work, we have investigated the process of electron-positron annihilation into a $\Sigma^{+}\bar{\Sigma}^{-}$ pair in the vicinity of the charmonium resonance $\psi(3770)$ for the energy of BESIII. Moreover, we also studied this process in the vicinity of new charmonium-like states in the same energy region, such as $\psi(4040)$, $\psi(4160)$, $Y(4230)$, $Y(4360)$, $\psi(4415)$, and $Y(4660)$. In the process $e^{+}e^{-}\to\Sigma^{+}+\bar{\Sigma}^{-}$, besides the Born mechanism, which is presented in the framework of QED, we also investigated two more contributions associated with the intermediate state of the charmonium $\psi(3770)$ and other new charmonium-like states, such as $\psi(4040)$, $\psi(4160)$, $Y(4230)$, $Y(4360)$, $\psi(4415)$, and $Y(4660)$. Thus, one of these contributions is a $D$ meson loop and the other is the contribution of a three-gluon mechanism. Here we should remind that the process $e^{+}e^{-}\to\Sigma^{+}+\bar{\Sigma}^{-}$ can also be initiated by the vector-charmonium state as $\psi(3770)$, as well as other charmonium-like states, which were mentioned above. We understand that since the photon, $\psi$ and other charmonium-like states are vector mesons, the structures of the corresponding cross-section distributions is therefore similar.

### v3

@@ IV The D-meson loop mechanism
$S_{D}(s)=\frac{\alpha g_{e}\beta G(s)}{48\pi^{2}s}\int\frac{dk}{i\pi^{2}}\frac{TrD(s,k^{2})}{(k^{2}-M_{\Xi}^{2})((k-q_{+})^{2}-M_{D}^{2})((k+q_{-})^{2}-M_{D}^{2})}\times$
@@
$Z_{D}(s)=\frac{1}{s}\int\frac{dk}{i\pi^{2}}\frac{TrD(s,k^{2})}{(k^{2}-M_{\Xi}^{2})((k-q_{+})^{2}-M_{D}^{2})((k+q_{-})^{2}-M_{D}^{2})}\times$
@@
where $TrD(s,k^{2})$ is the trace of the $\gamma$-matrices over the baryon line have the following form:
$TrD(s,k^{2})=Tr[(\hat{q_{+}}+M_{\Sigma})\gamma_{5}(\hat{k}+M_{\Xi})\gamma_{5}(\hat{q_{-}}-M_{\Sigma})(\hat{k}-M_{\Sigma})]=$
@@
Thus, using the Cutkosky cutting rules Cutkosky , we can replace the $D$-meson propagators in (31) through the $\delta$-functions, then we obtain the imaginary part of this quantity from $Z_{D}(s)$,
$2i\,\mbox{Im}\,Z_{D}(s)=\frac{(-2\pi i)^{2}}{s}\int\frac{dk}{i\pi^{2}}\frac{TrD(s,k^{2})}{k^{2}-M_{\Xi}^{2}}G_{\psi D\bar{D}}(s,(k+q_{-})^{2},(k-q_{+})^{2})\times$
@@
×$\quad\times TrD(s,-k_{(i)}^{2})\leavevmode\nobreak\ G_{\Xi D\Sigma}^{2}(-k_{(i)}^{2},M_{D}^{2}),\qquad s>4M_{D}^{2},$ (35)
@@
Now we want to consider the function $G_{\Xi D\Sigma}(k^{2},p^{2})$ from (31). Since $k^{2}<0$, in the $t$-channel the only dependence in the imaginary part of $Z_{D}$ is the off-mass shell of the $\Xi$ baryon.
We would like to note that in Ahmadov ; Bystritskiy ; BA we used a constant for the $\Lambda DP$-vertex based on the results of Reinders ; Navarra . But, in this paper we use for $\Xi D\Sigma$-vertex the constant that corresponds to the results of Reinders ; Choe .
@@ V The three-gluon mechanism
$Z_{3g}(s)=\frac{4}{\pi^{5}s}\int\frac{dk_{1}}{k_{1}^{2}}\frac{dk_{2}}{k_{2}^{2}}\frac{dk_{3}}{k_{3}^{2}}\frac{Tr3g\leavevmode\nobreak\ \delta(q-k_{1}-k_{2}-k_{3})}{((q_{+}-k_{1})^{2}-M_{\Sigma}^{2})((q_{-}-k_{3})^{2}-M_{\Sigma}^{2})},$ (44)
where $g_{col}$ is the color factor, $g_{col}(1/4)=\left<\Sigma|d^{ijk}\leavevmode\nobreak\ T^{i}T^{j}T^{k}|\Sigma\right>=15/2$. The $Tr3g$ in (44) is the product of traces over the lines of the $\Sigma$-hyperon and $c$-quark,
$Tr3g=\mbox{Tr}[\hat{Q}_{\alpha\beta\gamma}(\hat{k}_{c}+m_{c})\gamma^{\mu}(\hat{k}_{\bar{c}}-m_{c})]\cdot\mbox{Tr}\left[(\hat{q}_{+}+M_{\Sigma})\gamma^{\alpha}(\hat{q}_{+}-\hat{k}_{1}+M_{\Sigma})\gamma^{\beta}\right.\times$
@@
The factor $1/\sqrt{3}$ grants the correct normalization of this wave function.
@@ VI The Numerical results
For the process of $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ in the Born approximation, in Fig. 3 we show the dependence of the total cross as a function of the center-of-mass energy $\sqrt{s}$. It is seen that at the energy of $\sqrt{s}$ = 3.4 GeV, the total cross section has a maximum value and then, with increasing energy, the total cross section smoothly decreases. A comparison with the experimental data shows that our result agrees well with the experimental points. In Fig. 3, we constructed a plot according to the theoretical results for the total cross section in the Born approximation. Our theoretical result is compared with the experimental data from the BESIII.
We would like to note that the propagator between the $D^{0}\bar{D^{0}}$ vertices contains two types of hyperons – $\Sigma^{+}_{c},\Xi^{+}_{c}$, that is, there are different options. These hyperons are obtained from different combinations of quarks. When we performed the calculations, we took for the hyperon mass in the propagator the average value of the mass of the two hyperons.
It should be noted that the quantities $Z_{D}(s)$ from (31) and $Z_{3g}\left(s\right)$ from (44) are considered the main parts of the general cross section, which give the corresponding contributions ($D$-meson loop and three-gluons). In Fig. 6 we present the dependence of the quantity of $Z_{D}(s)$ on the total energy $\sqrt{s}$ in the range from the reaction threshold $\sqrt{s}=2M_{\Sigma}^{+}$ to $4.951\leavevmode\nobreak$GeV. It can be seen from the figure that in the real and imaginary parts of $Z_{D}(s)$ the quantity remains practically the same as in the case of the production of final states $p\bar{p}$ [Fig. 7(a) in Bystritskiy ], and $\Lambda\bar{\Lambda}$ [Fig. 6 in BA ], and $\Sigma^{0}\bar{\Sigma}^{0}$ [Fig. 6 in AAhmadov ]. Now in Fig. 7 we show the dependence of the real and imaginary parts of the quantity $Z_{3g}(s)$ on the total energy $\sqrt{s}$ in the range from the reaction threshold $\sqrt{s}=2M_{\Sigma}^{+}$ to $4.951\leavevmode\nobreak$GeV. We would like to note that in this work the general behavior of the curves of the quantity $Z_{3g}(s)$ by $\sqrt{s}$ is very different from the curves, which were produced in the final state $p\bar{p}$ [Fig.  7 (B) Bystritskiy ], $\Lambda\bar{\Lambda}$ [Fig. 7 BA ] and $\Sigma^{0}\bar{\Sigma}^{0}$ [Fig. 7 in AAhmadov ]. It can be seen from this graph that the characteristic large negative values of the quantity $Z_{3g}(s)$ remain the same, which gives a large relative phase with respect to the Born contribution to the amplitude. The resonance position $\psi(3770)$ in Figs. 6 and 7 is marked by a vertical dashed line.
@@
The contributions of the pure $D$-meson loop to the Born cross section are shown in Fig.8. In Fig.9 we illustrate the result of a pure three-gluon contribution. The theoretical results obtained for both of these contributions are compared with the data of the BESIII I26 collaboration.
@@
In Fig. 10 we present the results of the dependence of the total cross section in the energy range $\sqrt{s}$ = 3.4 - 5.0 GeV taking into account the $D$-loop and three-gluon mechanisms in the Born cross section and compare it with the BESIII I26 data. In this graph, we want to show that the peak is near the resonance of the $\psi(3770)$ charmonia and is clearly visible. It is necessary to note that in our model, we mainly consider the vicinity of $\psi(3770)$. But, in addition to this, we consider the vicinity with new charmoniumlike states in the same energy region, for example, $\psi(4040)$, $\psi(4160)$, $Y(4230)$, $Y(4360)$, $\psi(4415)$ and $Y(4660)$. Thus, we basically have the $\psi(3770)$ charmonium in the intermediate state, but in addition we calculate the total cross section taking into account the new charmoniumlike states in the intermediate state.
@@
Now in Fig. 11 we illustrate the dependence of the total relative phase $\phi_{\psi}$ from the center of mass energy $\sqrt{s}$. We determine the total relative phase $\phi_{\psi}$ as the contribution of the charmonium $\mathcal{M}_{\psi}$ to the amplitude with respect to the Born contribution of $\mathcal{M}_{B}$ without taking into account the Breit-Wigner factor, that is,
@@
In this present work, in addition to the contribution of the $\psi(3770)$ charmonium in the intermediate state, we also considered the contributions from charmoniumlike states, $\psi(4040)$,   $\psi(4160)$,   $Y(4230)$,  $Y(4360)$,  $\psi(4415)$, and $Y(4660)$ in the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ process. In Fig. 12 we present the total cross section at the center-of-mass of energy in the region $\sqrt{s}$ = 3.510 - 4.951 GeV according to the results of the contributions of charmonium (like) states, $\psi(4040)$,   $\psi(4160)$,   $Y(4230)$,  $Y(4360)$,  $\psi(4415)$, and $Y(4660)$ in the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ process, respectively. We would like to note that in Fig. 12 it is seen that the experimental result at the point for each resonance of charmonia $\psi(4040)$,   $\psi(4160)$,   $Y(4230)$,  $Y(4360)$,  $\psi(4415)$, and $Y(4660)$ are in very good agreement with our theoretical result.
@@ VII Summary and Conclusion
In the present work, we have investigated the process of electron-positron annihilation into a $\Sigma^{+}\bar{\Sigma}^{-}$ pair in the vicinity of the charmonium resonance $\psi(3770)$ for the energy of BESIII. Moreover, we also studied this process in the vicinity of new charmoniumlike states in the same energy region, such as $\psi(4040)$, $\psi(4160)$, $Y(4230)$, $Y(4360)$, $\psi(4415)$, and $Y(4660)$. In the process $e^{+}+e^{-}\to\Sigma^{+}+\bar{\Sigma}^{-}$, besides the Born mechanism, which is presented in the framework of QED, we also investigated two more contributions associated with the intermediate state of the charmonium $\psi(3770)$ and other new charmoniumlike states, such as $\psi(4040)$, $\psi(4160)$, $Y(4230)$, $Y(4360)$, $\psi(4415)$, and $Y(4660)$. Thus, one of these contributions is a $D$ meson loop and the other is the contribution of a three-gluon mechanism. Here we should remind that the process $e^{+}+e^{-}\to\Sigma^{+}+\bar{\Sigma}^{-}$ can also be initiated by the vector-charmonium state as $\psi(3770)$, as well as other charmoniumlike states, which were mentioned above. We understand that since the photon, $\psi$ and other charmoniumlike states are vector mesons, the structures of the corresponding cross-section distributions is therefore similar.